MICQ(1) MICQ(1) mICQ−,ICQ, micq[-v[]][-f] NOTE-THISFILEHASNOTYETBEENUPDATEDFORmICQ0.4.10. mICQICQ,ICQ,Mirabilis.mICQ,,,. -v[]. -frc. -l|-,-. -?,-h. mICQLinux,DOS,OS/2,Win32,BeOS,AmigaOS,GNUHURD-UNIX.mICQ 0.4.9.2Linux,AIX,HPUX,Tru64,AmigaOS,BeOSUNIX-.,,,-. 161998 MatthewD.Smithmicq.25199812. 052001 MatthewD.Smith-.Matthew,mICQBSDmICQ. 102002 0.4.9GPL,ICQ8.-Ru¨ digerKuhlmann. mICQ:micq-list@micq.org.http://www.micq.org/pipermail/micq-list/. HTTPhttp://www.micq.org/download. ,CVS, cvs-d:pserver:anonymous@micq.org:/usr/local/cvsrootcomicq micqrc(5),micq(7) ˜ /.micqrc mICQ--. ˜ /.micq/micqrc mICQ. PREFIX/share/micq/*.i18n ,mICQ. ˜ /.micq/i18n/*.i18n ˜ /.micq/history/ ,,,-. UK mICQ 1 MICQ(1) MICQ(1) Ru¨ digerKuhlmann<micq@ruediger-kuhlmann.de>(). JamesMorrison<ja2morrison@student.math.uwaterloo.ca>,JordiMallach<jordi@sindominio.net>DebianGNU/Linux.ArtemSokolov<sokolov@ukr.net>. UK mICQ 2